# Chirped DFT-s-OFDM: A new single-carrier waveform with enhanced LMMSE noise suppression

### v1

@@ Abstract
In this correspondence, a new single-carrier waveform, called chirped discrete Fourier transform spread orthogonal frequency division multiplexing (DFT-s-OFDM), is proposed for the sixth generation of communications. By chirping DFT-s-OFDM in the time domain, the proposed waveform maintains the low peak-to-average-power ratio (PAPR) of DFT-s-OFDM. Thanks to full-band transmission and symbols retransmission enabled by chirping and discrete Fourier transform (DFT) precoding, the proposed waveform can enhance noise suppression of linear minimum mean square error equalization. Its bit error rate (BER) upper bound and diversity order are derived using pairwise error probability. Simulation results confirm that the proposed waveform outperforms the state-of-the-art waveforms in terms of BER, output signal-to-noise-ratio, and PAPR.
@@ I Introduction
Transmission waveform has been regarded as one of important components in every generation of communications, enabling increased capacity and data rates, improved spectral and power efficiency, new applications, etc. Orthogonal frequency division multiplexing (OFDM) was adopted for downlink transmissions in the fourth (4G) and fifth generations (5G) of communications. For the sake of low peak-to-average-power-ratio (PAPR), discrete Fourier transform spread OFDM (DFT-s-OFDM) [1] was selected for uplink transmissions. However, they may be not suitable for high-mobility communications in the sixth generation (6G) of communications. 6G waveform design has thus attracted significant attention from both academia and industry recently.
The 6G waveform candidates in the literature can be classified into three categories: multi-carrier, single-carrier, and multi-carrier chirping. Multi-carrier category includes orthogonal time frequency space (OTFS) [2], orthogonal delay-Doppler division multiplexing (ODDM) [3], frequency-modulated OFDM (FM-OFDM) [4], etc. However, similar to OFDM, OTFS and ODDM based on multi-carrier have high PAPR, making it unsuited for low-power devices. Though FM-OFDM could reduce PAPR, its frequency spectrum utilization with lots of inactive subcarriers is inefficient. Meanwhile, the performance advantages of FM-OFDM over OTFS are not identified. Single-carrier category consists of orthogonal time sequency multiplexing (OTSM), whose PAPR is however as high as that of OTFS [5]. Orthogonal chirping division multiplexing (OCDM) [6] and affine frequency division multiplexing (AFDM) [7] belong to multi-carrier chirping category. AFDM is an enhanced version of OCDM to cater for high-mobility communications. However, they are based on multi-carrier and have high PAPR as well.
In this paper, a new single-carrier waveform called chirped DFT-s-OFDM is proposed for 6G communications. The contributions of this paper are summarized as follows:
@@
  • To the best of the authors’ knowledge, this is the first single-carrier chirping waveform candidate featuring low PAPR for 6G communications. In contrast, the existing 6G waveform candidates are multi-carrier, either without [2, 3, 4] or with chirping [6, 7], all suffering high PAPR.
@@
  • By using chirping and discrete Fourier transform (DFT) precoding, the proposed waveform enables full-band transmission and symbols retransmission, contributing to enhanced noise suppression of linear minimum mean square error (LMMSE) equalization. Simulation results confirm that it outperforms the existing waveforms, e.g., DFT-s-OFDM, AFDM, OTFS, and OFDM, in terms of bit error rate (BER) and output signal-to-noise-ratio (SNR).
  • Like multi-carrier chirping waveform, e.g., AFDM, the proposed single-carrier chirping waveform can also achieve full frequency diversity. Its BER upper bound and diversity order are derived based on pairwise error probability (PEP). Simulation results verify that the simulated BER approaches the derived BER upper bound.
@@
Figure 1: Block diagram of proposed chirped DFT-s-OFDM.
@@ II-A Modulation and Demodulation
The block diagram of proposed chirped DFT-s-OFDM is illustrated in Fig. 1. At transmitter side, the proposed chirped DFT-s-OFDM includes two main blocks, with DFT-s-OFDM modulation followed by chirping. The un-modulated time-domain data vector of length $M$ is denoted as $\mathbf{x}=[x[0],x[1],\cdots,x[M-1]]^{T}$. After $M$-point DFT, the frequency-domain data vector of length $M$ is given by $\mathbf{X}=\mathbf{F}_{M}\mathbf{x}$, with $\mathbf{F}_{M}$ being $M$-point DFT matrix. After subcarrier mapping, the data vector of length $N$ is written as $\mathbf{S}=\mathbf{P}\mathbf{X}$, with $\mathbf{P}$ being a subcarrier mapping matrix of size $N\times M$ ($N>M$). Interleaved subcarrier mapping with $\mathbf{P}=\mathbf{I}_{M}\otimes[1;\mathbf{0}_{(N/M-1)\times 1}]$ is used in this paper, with $\mathbf{I}_{M}$ being an identity matrix of size $M\times M$. The DFT spreading factor is defined as $\mathrm{SF}=N/M$. Inverse fast Fourier transform (FFT) is then implemented, resulting in time-domain signal $\mathbf{s}=\mathbf{F}_{N}^{H}\mathbf{S}$, with $\mathbf{F}_{N}$ being $N$-point FFT matrix. The DFT-s-OFDM signal $\mathbf{s}$ is then chirped with a linear chirp signal $\mathbf{c}$ of chirp rate $c$, obtaining the chirped DFT-s-OFDM signal denoted as $\mathbf{s}_{\mathrm{c}}$:
@@
with $\mathbf{c}=[1,e^{j\pi c},\cdots,e^{j\pi cn^{2}},\cdots,e^{j\pi c(N-1)^{2}}]$.
A cyclic prefix of length $L_{\mathrm{cp}}$ ($L_{\mathrm{cp}}\geq L$) is appended to $\mathbf{s}_{\mathrm{c}}$ before transmission, with $L$ being the channel order. $h[n,l]$ is defined as channel gain of $l$-th channel path at $n$-th time instant, with $l=0,1,\cdots,L$ and $n=0,1,\cdots,N-1$. Define $\bar{f}_{\textrm{max}}=\frac{f_{\textrm{c}}v}{c}/\Delta f$ as the normalized maximum Doppler frequency with respect to subcarrier spacing, with $f_{\textrm{c}}$, $v$, $c$, and $\Delta f$ being carrier frequency, velocity, speed of light, and subcarrier spacing, respectively. $h[n,l]$ is given by
$h[n,l]=\sum\nolimits_{p=0}^{L}h_{p}e^{j2\pi v_{p}n/N}\delta(l-p),$ (2)
with $h_{p}$ and $v_{p}$ ($v_{p}\in[-\bar{f}_{\textrm{max}},\bar{f}_{\textrm{max}}]$) being the channel gain and normalized Doppler shift of $p$-th path. The time-domain circulant channel matrix is expressed as
$\mathbf{H}_{\rm{t}}=\sum\nolimits_{l=0}^{L}h_{l}\mathbf{D}_{l}\bm{\Pi}^{l},$ (3)
where $\mathbf{D}_{l}=\mathrm{diag}\{1,\cdots,e^{j2\pi v_{l}n/N},\cdots,e^{j2\pi v_{l}(N-1)/N}\}$ and $\bm{\Pi}$ is the forward cyclic-shift matrix defined as (25) of [7].
Figure 2: Example of time-frequency diagrams, with $N=8$ and $c=1/8$.
@@
OFDM channel estimation and equalization schemes can then be exploited to recover the frequency-domain symbols. The transmitted time-domain data symbols $\mathbf{x}$ can be recovered by discarding unused subcarriers and performing inverse $M$-point DFT of the estimated frequency-domain symbols.
@@ II-B Time-Frequency Diagram
The time-frequency diagrams of DFT-s-OFDM and proposed chirped DFT-s-OFDM are illustrated in Fig. 2, with $N=8$ and $c=1/8$. Regarding DFT-s-OFDM with $\rm{SF}=2$, information symbols are mapped in the time domain in four different colors and then are repeated in the time domain due to DFT precoding, with each symbol occupying the assigned subcarrier bands. The un-assigned subcarrier bands are indicated in white color and can be used by other users for data transmission. A linear chirp signal with chirp rate $c=1/8$ is used, implying that the frequency is increased by one subcarrier spacing per time instant. After chirping, the data symbols, e.g., in blue and orange color, can hop to the un-assigned subcarrier bands. Thus, the chirped DFT-s-OFDM enables data transmission over the full band, leading to enhanced noise suppression in the frequency domain. Besides, when $\rm{SF}$ increases to $4$ and each symbol would be repeated $4$ times in the time domain, additional noise suppression can be obtained. Fig. 7 in Section IV will confirm that the increase of $\rm{SF}$ would reduce output noise power and BER. For the sake of full band transmission, the chirp rate should be properly selected by using $c=(a+b\times\rm{SF})/N$, with $a=1,2,\cdots,\rm{SF}-1$ and $b$ being zero or positive integer value.
It is noteworthy that chirp signal was applied with DFT-s-OFDM in [1], which however is used in the frequency domain and at transmitter side only. It also presents limited performance gains over DFT-s-OFDM, especially when linear chirp is used and DFT spreading factor is big [1]. In contrast, chirp signal is used in the time domain at both transmitter side and receiver side for the proposed waveform, referred to as chirping and dechirping in Fig. 1. According to [7], chirping in the time domain would have a greater impact than that in the frequency domain. As will be seen in Section IV, the proposed chirped DFT-s-OFDM with chirping in the time domain would have superior BER and output SNR over DFT-s-OFDM, especially when it has a big DFT spreading factor.
@@ III Performance Analysis
The error performance of proposed chirped DFT-s-OFDM is investigated using PEP analysis of maximum likelihood (ML) equalizer. Then, the output SNR of proposed waveform is derived and analyzed using LMMSE equalizer.
@@ III-A Error Performance Analysis Using ML Equalizer
The system model in (5) can be rewritten as
$\mathbf{y}$ $=\sum\nolimits_{l=0}^{L}\mathbf{F}_{N}\mathbf{C}^{H}\mathbf{D}_{l}\bm{\Pi}^{l}\mathbf{C}\mathbf{F}_{N}^{H}\mathbf{P}\mathbf{F}_{M}\mathbf{x}h_{l}+\mathbf{z}.$ (6)
Denote $\mathbf{E}_{l}(\mathbf{x})=\mathbf{F}_{N}\mathbf{C}^{H}\mathbf{D}_{l}\bm{\Pi}^{l}\mathbf{C}\mathbf{F}_{N}^{H}\mathbf{P}\mathbf{F}_{M}\mathbf{x}$. The received frequency-domain signal vector $\mathbf{y}$ is reexpressed as
@@
with $\mathbf{E}=[\mathbf{E}_{0},\mathbf{E}_{1},\cdots,\mathbf{E}_{L}]$ and $\mathbf{h}=[h_{0},h_{1},\cdots,h_{L}]^{T}$.
Denote $\{\mathbf{x}\rightarrow\hat{\mathbf{x}}\}$ as the pairwise error event, where $\mathbf{x}$ represents the transmitted signal vector and $\hat{\mathbf{x}}$ the erroneously detected signal vector using ML equalizer. $\mathbf{x}$ and $\hat{\mathbf{x}}$ are chosen from a certain modulation alphabet, i.e., $\mathbf{x}\in\mathbb{K}$ and $\hat{\mathbf{x}}\in\mathbb{K}$, but with $\mathbf{x}\neq\hat{\mathbf{x}}$. Define $\bm{\Theta}(\mathbf{x},\hat{\mathbf{x}})=(\mathbf{E}(\mathbf{x})-\mathbf{E}(\hat{\mathbf{x}}))^{H}(\mathbf{E}(\mathbf{x})-\mathbf{E}(\hat{\mathbf{x}}))$. The rank of $\bm{\Theta}(\mathbf{x},\hat{\mathbf{x}})$ and its non-zero eigenvalues are denoted as $R$ and $\{\lambda_{1},\lambda_{2},\cdots,\lambda_{R}\}$. Assume each entry in $\mathbf{h}$ is an independent identically distributed complex Gaussian random variable, i.e., $\mathbf{h}\sim\mathcal{CN}(0,\frac{1}{L+1}\mathbf{I}_{L+1})$. The SNR for each data symbol is denoted as $\gamma$, i.e., $\gamma=\frac{1}{\sigma^{2}}$. Following the derivation in [8], the PEP can be expressed as
$P_{E}(\mathbf{x}\rightarrow\hat{\mathbf{x}})\approx\frac{1}{12}\prod\nolimits_{r=1}^{R}\frac{1}{1+\frac{\gamma\lambda_{r}}{4(L+1)}}+\frac{1}{4}\prod\nolimits_{r=1}^{R}\frac{1}{1+\frac{\gamma\lambda_{r}}{3(L+1)}}.$ (8)
@@
$P_{E}(\mathbf{x}\rightarrow\hat{\mathbf{x}})\approx\frac{1}{12}\left[\left(\prod\nolimits_{r=1}^{R}\lambda_{r}\right)^{\frac{1}{R}}\frac{\gamma}{4(L+1)}\right]^{-R}+$
$\frac{1}{4}\left[\left(\prod\nolimits_{r=1}^{R}\lambda_{r}\right)^{\frac{1}{R}}\frac{\gamma}{3(L+1)}\right]^{-R}.$ (9)
The average BER can then be obtained by using the union bound technique [8] as follows:
@@
By calculating $G_{D}$ using (11), it is found to be equal to the number of channel paths.
Fig. 3 shows the derived BER upper bound (10) and simulated BER of proposed chirped DFT-s-OFDM using ML equalizer. The DFT size and IFFT size of chirped DFT-s-OFDM are $M=2$ and $N=8$. The maximum Doppler frequency, velocity, and subcarrier spacing are $f_{\rm{max}}=2\,\rm{KHz}$, $v=500\,\rm{km/h}$, and $\Delta f=15\,\rm{KHz}$. Binary phase shift keying (BPSK) modulation is adopted. According to Fig. 3, the simulated BER is close to the derived BER upper bound especially at high SNRs. The slope of BER curves corresponds to the diversity order of chirped DFT-s-OFDM, which is found to be the number of channel paths and is consistent with (11).
Figure 3: Derived BER upper bound (10) and simulated BER of proposed chirped DFT-s-OFDM using ML equalizer.
(a) Output signal power ($\rm{Tr}\{\mathbf{H}_{\rm{eff}}^{H}\mathbf{G}^{H}\mathbf{G}\mathbf{H}_{\rm{eff}}\}$).
(b) Output noise power $(\rm{Tr}\{\sigma^{2}\mathbf{G}^{H}\mathbf{G}\})$.
(c) Output SNR (the ratio of Figs. 4a and 4b).

### v2

@@ Abstract
In this paper, a new single-carrier waveform, called chirped discrete Fourier transform spread orthogonal frequency division multiplexing (DFT-s-OFDM), is proposed for the sixth generation of communications. As the initial study on this waveform, its performance is analyzed and evaluated in single-user uplink multiple access communications, using a delay-Doppler channel model with independent and identically distributed path amplitudes and different integer delays. By chirping DFT-s-OFDM in the time domain, it maintains the low peak-to-average-power ratio of DFT-s-OFDM. Thanks to full-band transmission and symbols retransmission enabled by chirping and discrete Fourier transform precoding, it enhances noise suppression of linear minimum mean square error equalization. By using pairwise error probability analysis, the derived bit error rate upper bound is close to the simulated ones and the diversity order analysis confirms that it achieves full frequency diversity.
@@ I Introduction
Transmission waveform has been regarded as one of important components in every generation of communications. Orthogonal frequency division multiplexing (OFDM) was adopted for downlink transmissions in the fourth (4G) and fifth generations (5G) of communications. For the sake of low peak-to-average-power-ratio (PAPR), discrete Fourier transform spread OFDM (DFT-s-OFDM) [1, 2, 3, 4] was selected for uplink transmissions. However, they may not be well-suited for the high-mobility applications in the sixth generation (6G) of communications. 6G waveform design has thus attracted significant attention from both academia and industry recently.
The 6G waveform candidates in the literature include orthogonal time frequency space (OTFS) [5, 6, 7], orthogonal delay-Doppler division multiplexing (ODDM) [8], orthogonal time sequency multiplexing (OTSM) [9], frequency-modulated OFDM (FM-OFDM) [10], affine frequency division multiplexing (AFDM) [11], etc. However, OTFS, ODDM and OTSM all have high PAPR. PAPR reduction [12] has been studied for OTFS. However, it introduces clipping noise and additional compensation methods are required [13]. FM-OFDM is a variant of constant-envelope OFDM (CE-OFDM) [14] and inherits its perfect PAPR. It performs phase accumulation of OFDM symbols before phase modulator and exhibits higher resilience to Doppler frequency than CE-OFDM [10]. However, its benefits come at the expense of expanding bandwidth. Similarly to orthogonal chirping division multiplexing [15], AFDM adds chirping in the frequency and time domain before and after OFDM modulation, respectively. To cater for high-mobility communications, its chirp rates are properly selected according to maximum Doppler frequency. However, AFDM, as a variation of OFDM, retains its high PAPR.
In this correspondence, a new single-carrier waveform, called chirped DFT-s-OFDM, is proposed for 6G communications, with an initial study conducted in single-user uplink multiple access communications (MAC) using the delay-Doppler channel model [11] with independent and identically distributed (i.i.d.) path amplitudes and different integer delays. The contributions of this paper are summarized as follow:
@@
  • To the best of the authors’ knowledge, this is the first single-carrier chirping waveform featuring low PAPR for 6G communications. In contrast, the existing 6G waveform candidates, including OTFS [5, 6, 7], ODDM [8], OTSM [9], and AFDM [11], suffer high PAPR.
@@
  • By using chirping and discrete Fourier transform (DFT) precoding, the proposed waveform enables full-band transmission and symbols retransmission, leading to enhanced noise suppression of linear minimum mean square error (LMMSE) equalization. Simulation results confirm that it outperforms DFT-s-OFDM, AFDM, OTFS, and OFDM in terms of bit error rate (BER) and output signal-to-noise-ratio (SNR) in single-user uplink MAC.
  • The BER upper bound and diversity order of proposed waveform are derived based on pairwise error probability (PEP). The simulated BER closely approaches the derived BER upper bound, and the derived diversity order corresponds to the number of integer delays, confirming full frequency diversity over delay-Doppler channels with i.i.d. path amplitudes and different integer delays.
@@
Figure 1: a) Block diagram of proposed chirped DFT-s-OFDM, and b) time-frequency diagrams of (chirped) DFT-s-OFDM, with $N=8$ and $c_{\rm{r}}=1/8$.
@@ II-A Modulation and Demodulation
Considering single-user uplink MAC, the block diagram of proposed chirped DFT-s-OFDM is illustrated in Fig. 1a. At transmitter side, it includes two main blocks, with DFT-s-OFDM modulation followed by chirping. The un-modulated time-domain data vector of length $M$ is denoted as $\mathbf{x}=[x[0],x[1],\cdots,x[M-1]]^{T}$. The elements of data vector should be i.i.d., i.e., $\mathbb{E}\{\mathbf{x}\mathbf{x}^{H}\}=\mathbf{I}_{M}$, where each data symbol has unit power and $\mathbf{I}_{M}$ is an identity matrix of size $M$. After $M$-point DFT, the frequency-domain data vector of length $M$ is given by $\mathbf{X}=\mathbf{F}_{M}\mathbf{x}$, with $\mathbf{F}_{M}$ being $M$-point DFT matrix. After subcarrier mapping, the data vector of length $N$ is written as $\mathbf{S}=\mathbf{P}\mathbf{X}$, with $\mathbf{P}$ being a subcarrier mapping matrix of size $N\times M$ ($N>M$). The DFT spreading factor is defined as $\mathrm{SF}=N/M$ with $\mathrm{SF}>1$. For the sake of low PAPR, interleaved subcarrier mapping with an integer value of $\mathrm{SF}$ [2, 3, 4] is considered in this paper, yielding $\mathbf{P}=\mathbf{I}_{M}\otimes[1;\mathbf{0}_{(\mathrm{SF}-1)\times 1}]$. Inverse fast Fourier transform (FFT) is then implemented, resulting in time-domain signal $\mathbf{s}=\mathbf{F}_{N}^{H}\mathbf{S}$, with $\mathbf{F}_{N}$ being $N$-point FFT matrix. The DFT-s-OFDM signal $\mathbf{s}$ is then chirped with a linear chirp signal $\mathbf{c}$ of chirp rate $c_{\rm{r}}$, obtaining the chirped DFT-s-OFDM signal denoted as $\mathbf{s}_{\mathrm{c}}$ as follow:
@@
with $\mathbf{c}=[1,e^{j\pi c_{\rm{r}}},\cdots,e^{j\pi c_{\rm{r}}n^{2}},\cdots,e^{j\pi c_{\rm{r}}(N-1)^{2}}]$.
A cyclic prefix of length $L_{\mathrm{cp}}$ ($L_{\mathrm{cp}}\geq L-1$) is appended to $\mathbf{s}_{\mathrm{c}}$ before transmission, with $L$ being the number of resolvable paths. $h[n,l]$ is defined as channel gain of $l$-th channel path at $n$-th time instant, with $l=1,2,\cdots,L$ and $n=0,1,\cdots,N-1$. Define $\bar{f}_{\textrm{max}}=\frac{f_{\textrm{c}}v}{c}/\Delta f$ as the normalized maximum Doppler frequency with respect to subcarrier spacing, with $f_{\textrm{c}}$, $v$, $c$, and $\Delta f$ being carrier frequency, velocity, speed of light, and subcarrier spacing, respectively. The delay-Doppler channel model in [11] is used in this paper:
$h[n,l]=\sum\nolimits_{p=1}^{L}h_{p}e^{j2\pi\nu_{p}n/N}\delta(l-l_{p}),$ (2)
with $h_{p}$, $\nu_{p}$, and $l_{p}$ being the channel amplitude, normalized Doppler shift, and integer delay of $p$-th path. $h_{p}$ is modeled as an i.i.d. complex Gaussian random variable with zero mean and variance of $1/L$ and $\nu_{p}$ is randomly selected from $-\bar{f}_{\textrm{max}}$ to $\bar{f}_{\textrm{max}}$ [11]. This paper assumes that different paths have different integer delays. However, (2) can be extended to the scenarios with multiple paths sharing the identical integer delay by setting $l_{p}=l_{q}$ for $p$-th and $q$-th path, while with $h_{p}\neq h_{q}$ and $\nu_{p}\neq\nu_{q}$ [11]. The time-domain channel matrix is expressed as $\mathbf{H}_{\rm{t}}=\sum\nolimits_{l=1}^{L}h_{l}\mathbf{D}_{l}\bm{\Pi}^{l}$, where $\mathbf{D}_{l}=\mathrm{diag}\{1,\cdots,e^{j2\pi\nu_{l}n/N},\cdots,e^{j2\pi\nu_{l}(N-1)/N}\}$ and $\bm{\Pi}$ is the forward cyclic-shift matrix defined as (25) of [11].
@@ II-B Time-Frequency Diagram
The time-frequency diagrams of (chirped) DFT-s-OFDM are illustrated in Fig. 1b for a certain user, with $N=8$ and $c_{\rm{r}}=1/8$. Regarding DFT-s-OFDM with $\rm{SF}=2$, information symbols are mapped in the time domain in four different colors and then are repeated in the time domain with reduced amplitude, with each symbol occupying the assigned subcarrier bands [2, 3, 4]. The un-assigned subcarrier bands are indicated in white color and can be used by other users. A linear chirp signal with chirp rate $c_{\rm{r}}=1/N$ is used, implying that the subcarrier frequency is increased by one subcarrier spacing per time instant. The original subcarrier $k$ ($k=0,1,\cdots,N-1$) at $n=0$ would be changed to $k+n$ at time instant $n$. For example, the subcarrier 0 remains at subcarrier 00 at $n=0$, hops to subcarrier $1$ at $n=1$, etc. After chirping, the data symbols, e.g., in blue and orange color, can hop to the un-assigned subcarrier bands to enable full band transmission. It then provides more observations than unknowns, and the redundant information from more equations can thus enhance noise suppression in the frequency domain. Besides, when $\rm{SF}$ increases to $4$ and each symbol would be repeated $4$ times in the time domain, additional noise suppression can be obtained.
Note that chirp signal was applied with DFT-s-OFDM in [1], which however is used in the frequency domain and at transmitter only. It also presents limited performance gains over DFT-s-OFDM [1]. In contrast, chirp signal is used in the time domain at both transmitter and receiver for the proposed waveform, referred to as chirping and dechirping in Fig. 1a. According to [11], chirping in the time domain is expected to have a greater impact than that in the frequency domain.
@@ III-A Error Performance Analysis Using ML Equalizer
This subsection aims to apply PEP analysis technique of maximum likelihood (ML) equalizer to derive the BER upper bound and diversity order of proposed chirped DFT-s-OFDM. The system model in (4) is rewritten as
$\mathbf{y}=\sum\nolimits_{l=1}^{L}\mathbf{F}_{N}\mathbf{C}^{H}\mathbf{D}_{l}\bm{\Pi}^{l}\mathbf{C}\mathbf{F}_{N}^{H}\mathbf{P}\mathbf{F}_{M}\mathbf{x}h_{l}+\mathbf{z}.$ (5)
Denote $\mathbf{E}_{l}(\mathbf{x})=\mathbf{F}_{N}\mathbf{C}^{H}\mathbf{D}_{l}\bm{\Pi}^{l}\mathbf{C}\mathbf{F}_{N}^{H}\mathbf{P}\mathbf{F}_{M}\mathbf{x}$. $\mathbf{y}$ is expressed as
@@
with $\mathbf{E}=[\mathbf{E}_{1},\mathbf{E}_{2},\cdots,\mathbf{E}_{L}]$ and $\mathbf{h}=[h_{1},h_{2},\cdots,h_{L}]^{T}$. $\mathbf{x}$ can be estimated by using ML equalizer:
$\mathbf{\hat{x}}=\min_{\mathbf{\tilde{x}}\in\mathbb{K}}\lVert\mathbf{y}-\mathbf{E}(\mathbf{\tilde{x}})\mathbf{h}\rVert,$ (7)
where $\mathbf{\tilde{x}}$ denotes the sequence candidate of data symbols and each element is chosen from a certain modulation alphabet $\mathbb{K}$. ML equalizer would incur high complexity, making it less suitable for big data vector size ($M$) and big symbol alphabet size ($Q$). Nevertheless, persurvivor processing (PSP) can be used to approximate ML algorithm with low complexity [16].
Denote $\{\mathbf{x}\rightarrow\hat{\mathbf{x}}\}$ as the pairwise error event, where $\mathbf{x}$ represents the transmitted signal vector and $\hat{\mathbf{x}}$ the erroneously detected signal vector using ML equalizer. $\mathbf{x}$ and $\hat{\mathbf{x}}$ are chosen from a certain modulation alphabet, i.e., $\mathbf{x}\in\mathbb{K}$ and $\hat{\mathbf{x}}\in\mathbb{K}$, but with $\mathbf{x}\neq\hat{\mathbf{x}}$. Define $\bm{\Theta}(\mathbf{x},\hat{\mathbf{x}})=(\mathbf{E}(\mathbf{x})-\mathbf{E}(\hat{\mathbf{x}}))^{H}(\mathbf{E}(\mathbf{x})-\mathbf{E}(\hat{\mathbf{x}}))$. The rank of $\bm{\Theta}(\mathbf{x},\hat{\mathbf{x}})$ and its non-zero eigenvalues are denoted as $R$ and $\{\lambda_{1},\lambda_{2},\cdots,\lambda_{R}\}$. The SNR for each data symbol is denoted as $\gamma$, i.e., $\gamma=\frac{1}{\sigma^{2}}$. Following the derivation in [6] and [7], the PEP can be expressed as
$P_{E}(\mathbf{x}\rightarrow\hat{\mathbf{x}})\approx\frac{1}{12}\prod\nolimits_{r=1}^{R}\frac{1}{1+\frac{\gamma\lambda_{r}}{4L}}+\frac{1}{4}\prod\nolimits_{r=1}^{R}\frac{1}{1+\frac{\gamma\lambda_{r}}{3L}}.$ (8)
@@
$P_{E}(\mathbf{x}\rightarrow\hat{\mathbf{x}})\approx\frac{1}{12}\left[\left(\prod\nolimits_{r=1}^{R}\lambda_{r}\right)^{\frac{1}{R}}\frac{\gamma}{4L}\right]^{-R}+$
$\frac{1}{4}\left[\left(\prod\nolimits_{r=1}^{R}\lambda_{r}\right)^{\frac{1}{R}}\frac{\gamma}{3L}\right]^{-R}.$ (9)
The average BER can then be obtained by using the union bound technique [6, 7] as follow:
@@
With $L=3$, Table I shows the diversity order $G_{D}$ of (chirped) DFT-s-OFDM computed using (11) with different combinations of Doppler shifts. The diversity order for DFT-s-OFDM is calculated by substituting $c_{\rm{r}}=0$ to (11). It can be seen that DFT-s-OFDM cannot achieve full frequency diversity and its diversity order varies with Doppler shifts. In contrast, the proposed waveform can exploit full frequency diversity and shows high resilience to Doppler shifts. To enable full frequency diversity, the chirp rate should be chosen from $c_{\rm{r}}=(a+b\times\mathrm{SF})/N$ to enable full bandwidth transmission, with $a=1,3,\cdots,\rm{SF}-1$ and $b$ being zero or positive integer value. For example, with $\mathrm{SF}=4$, $c_{\rm{r}}$ should be chosen from $\{1/N,3/N,5/N,...\}$.
TABLE I: Diversity order $G_{D}$ of (chirped) DFT-s-OFDM with $L=3$ calculated using (11). Note that for D: $v_{0}$, $v_{1}$, and $v_{2}$ are randomly generated from $-\bar{f}_{\textrm{max}}$ to $\bar{f}_{\textrm{max}}$ and the diversity order is calculated by averaging over $100,000$ realizations.
Doppler shifts DFT-s-OFDM Proposed
$[v_{0},v_{1},v_{2}]$ chirped DFT-s-OFDM
A: $[0,0,0]$ $1$ $3$
B: $[-\bar{f}_{\textrm{max}},0,\bar{f}_{\textrm{max}}]$ $2$ $3$
C: $[-\bar{f}_{\textrm{max}},-\bar{f}_{\textrm{max}},\bar{f}_{\textrm{max}}]$ $2$ $3$
D: Random $1.9$ $3$
Figure 2: Derived BER upper bound (10) and simulated BER of (a) (chirped) DFT-s-OFDM with $L=3$, and (b) proposed chirped DFT-s-OFDM.
Fig. 2 shows the derived BER upper bound (10) and simulated BER of (chirped) DFT-s-OFDM using ML equalizer. The DFT size and IFFT size are $M=2$ and $N=8$. The maximum Doppler frequency, velocity, and subcarrier spacing are $f_{\rm{max}}=2\,\rm{KHz}$, $v=500\,\rm{km/h}$, and $\Delta f=15\,\rm{KHz}$. Binary phase shift keying (BPSK) modulation is adopted. According to Fig. 2, the simulated BER is close to the derived BER upper bound especially at high SNRs. The BER curves for DFT-s-OFDM in Fig. 2a exhibit a flatter slope, suggesting that DFT-s-OFDM has a lower diversity order than chirped DFT-s-OFDM. For chirped DFT-s-OFDM, its diversity order calculated from its BER curvers is $L$ in Fig. 2b.
The above analysis of proposed waveform assumes all paths have different integer delays and i.i.d. path amplitudes. When some paths share the same integer delays, in addition to frequency diversity, Doppler/time diversity could be exploited thanks to Doppler spread [5] and a detailed analysis will be addressed in future work. If path amplitudes are correlated (not i.i.d.), full frequency diversity might not be attained.
(a) Output signal power ($\mathrm{diag}\{\mathbf{H}_{\rm{eff}}^{H}\mathbf{G}^{H}\mathbf{G}\mathbf{H}_{\rm{eff}}\}$).
(b) Output noise power $(\sigma^{2}\mathrm{diag}\{\mathbf{G}\mathbf{G}^{H})\}$.
(c) Output SNR (the ratio of Figs. 3a and 3b).